What acceleration will give the man with mass m = 75 kg, force F = 150N, when the angle it forms with the horizontal direction is 35º. It is known that cos 35º = 0.82. Other forces acting on the human feet in the horizontal direction, do not take them into account.


1
Expert's answer
2020-12-06T17:18:31-0500

The acceleration "a" in the horizontal direction is given by the Newton's second law:


"ma = F_{horizontal}"

where


"F_{horizontal} = F\cos35\degree = 150\cdot \cos35\degree = 150\cdot 0.82 = 123N"

Thus, obtain:


"a = \dfrac{F_{horizontal}}{m} = \dfrac{123N}{75kg} = 1.64m/s^2"

Answer. 1.64 m/s^2.